OpenStudy (anonymous):

What value of k solves the equation k^-4 = 1/256? A. –64 B. 8 C. 4 D. 64

OpenStudy (anonymous):

@math&ing001

OpenStudy (math&ing001):

k^-4 = 1/256 k^4 = 256 k^4 = 4^4 k=4 PS: k=-4 also works.